# Decentralized Parametric Stability Certificates for Grid-Forming Converter Control

### v5

@@ 1 Introduction
The transition to future power systems is characterized by a substantial increase in the share of power electronics (PE)-based generation devices. This shift introduces significant changes in system dynamics, where the interactions between PE-based generation and the remainder of the power network are not fully understood yet [1]. In particular, the fast and complex dynamics of PE-based devices can lead to unexpected effects (e.g., overvoltages [2], subsynchronous oscillations [3], etc.), posing considerable challenges for maintaining system stability. Consequently, there is a pressing need for a deeper understanding of how PE-based devices interact with one another or with the grid and for the development of robust stability frameworks to ensure stable operation.
@@
Our contribution improves significantly over prior works on stability certification in GFM converter systems. Unlike [7, 8], which are limited to single-input single-output (SISO) dynamics, our approach accommodates multiple-input multiple-output (MIMO) dynamics coupling frequency and voltage, thus providing a more comprehensive analysis of the system behavior. Additionally, we incorporate the network dynamics, extending beyond the quasi-stationary or zero-power flow approximations in [7, 9, 10]. Our decentralized stability conditions enable the use of heterogeneous GFM models, improving over the homogeneity assumptions in [8, 11]. Moreover, unlike [12], our approach does not require detailed knowledge of the network configuration. Finally, our stability conditions are independent of the operating point, provided small bus angle differences, unlike [9], where the the stability conditions depend on specific operating points. Collectively, these attributes allow our framework to provide, to the best of our knowledge, the most general, explicitly parametric, and practically applicable small-signal stability certificates for interconnected GFM devices in the literature. Since our stability conditions are readily actionable for control tuning and grid-code design, they can ensure a stable and scalable integration of PE-based generation in future power systems.
@@ 2.1 Preliminaries
Let $\mathbb{R}$ denote the set of real numbers, $\mathbb{N}$ the set of positive integers, and $\mathbb{C}$ the set of complex numbers with imaginary unit $\mathrm{j}$. We use $I_{n}$ to denote the $n$-by-$n$ identity matrix (abbreviated as $I$ when the dimensions can be inferred from the context). We use $\otimes$ to denote the Kronecker product. We use $A=\mathrm{diag}(A_{1},A_{2},...,A_{k}),\,k\in\mathbb{Z}$ to denote the block-diagonal matrix with blocks $A_{1},A_{2},...,A_{k}$. The Euclidean norm of a matrix $A$ is defined as $||A||_{F}=\sqrt{\textstyle\sum_{i,j}|a_{ij}|^{2}}$, where $a_{ij}$ is the entry in the $i$th row and $j$th column of $A$.
@@
To establish our main stability results in Section 4, we require the following lemma which follows from the Gershgorin’s Circle Theorem [13, Thm. 6.1.10]:
@@
Further, we review the recent concept of phases of complex matrices based on the matrix’s numerical range [14]. Namely, the numerical range of a complex matrix $A\in\mathbb{C}^{n\times n}$ is defined as $W(A)=\{x^{\star}Ax:x\in\mathbb{C}^{n},||x||=1\}$. If $0\notin W(A)$, then $A$ is said to be a sectorial matrix. For a sectorial $A$, there exists a nonsingular matrix $T$ and a diagonal unitary matrix $D$ such that $A=T^{\star}DT$ [14]. The diagonal elements of $D$ are distributed in an arc on the unit circle with length smaller than $\pi$. Then, the phases of $A$, denoted by
@@ 2.2 Transfer Functions & Stability
Let $u(t)\in\mathbb{R}^{n}$ be the input and $y(t)\in\mathbb{R}^{n}$ the output signal of a MIMO LTI system. The transfer matrix $H(s)\in\mathbb{C}^{n\times n}$ describes the input-output system response in the frequency domain as $y(s)=H(s)u(s)$, where $y(s)$ and $u(s)$ are the Laplace transformations of the output and input, respectively.
@@
$\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\underset{\eqqcolon\,H_{1}\#H_{2}(s)}{\underbrace{\begin{bmatrix}(I+H_{1}H_{2})^{-1}H_{1}&-(I+H_{1}H_{2})^{-1}H_{1}H_{2}\\ H_{2}(I+H_{1}H_{2})^{-1}H_{1}&H_{2}(I+H_{1}H_{2})^{-1}\end{bmatrix}}}\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}$ (2)
@@
If there are no RHP pole-zero cancellations between $H_{1}(s)$ and $H_{2}(s)$, then stability of one closed-loop transfer matrix implies stability of the others [15, Thm 4.7]:
@@ 2.3 Small-Phase & Passivity Theory
For LTI systems, the property of passivity is equivalent to the property of positive realness [14, 16, 17, 18].
Definition 2 (Passive Transfer Matrix [14]).
A $n\times n$ real rational proper transfer matrix $H(s)$ is passive if
@@
    any purely imaginary pole $\mathrm{j}\omega$ of any element of $H(s)$ is a simple pole and $\text{lim}_{s\rightarrow\mathrm{j}\omega}(s-\mathrm{j}\omega)H(s)$ is positive semi-definite Hermitian.
Definition 3 (Strictly Passive Transfer Matrix [14]).
A $n\times n$ real rational proper transfer matrix $H(s)$ is strictly passive if
@@
    $H(\mathrm{j}\omega)+H^{\star}(\mathrm{j}\omega)\succ 0$ for any $\omega\in[-\infty,\infty]$.
We now present a generalization of the passivity theorem [17, 18] for the special case of LTI systems as follows:
@@
Consider two LTI systems $H_{1}(s)$ and $H_{2}(s)$ in negative feedback configuration, as shown in Fig. 1(a). The feedback system is internally feedback stable if $H_{1}(s)$ is strictly passive and $H_{2}(s)$ is passive.
The above theorem is widely known in the SISO case [17], but we could not locate a proof in this general MIMO setting. We thus provide a self-contained proof below, relying on the Generalized Small-Phase Theorem [14, Thm 7.1]:
Lemma 3 (Generalized Small-Phase Theorem).
Let $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ be frequency-wise sectorial and $H_{2}(s)$ be semi-stable frequency-wise semi-sectorial with $\mathrm{j}\Omega$ being the set of poles on the imaginary axis. Then, $H_{1}\#H_{2}$ in Fig. 1(a) is internally feedback stable if
$\begin{split}&\underset{\omega\in[0,\infty]\backslash\Omega}{\mathrm{sup}}\left[\overline{\phi}(H_{1}(\mathrm{j}\omega))+\overline{\phi}(H_{2}(\mathrm{j}\omega))\right]<\pi\quad\text{and}\\ &\underset{\omega\in[0,\infty]\backslash\Omega}{\mathrm{inf}}\left[\underline{\phi}(H_{1}(\mathrm{j}\omega))+\underline{\phi}(H_{2}(\mathrm{j}\omega))\right]>-\pi.\end{split}$ (3)
@@
We prove Theorem 1 by applying Lemma 3 as follows: Since $H_{1}(s)$ is strictly passive, it follows that $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ is frequency-wise sectorial with phases $\left[\underline{\phi}(H_{1}(\mathrm{j}\omega)),\overline{\phi}(H_{1}(\mathrm{j}\omega))\right]\subset(-\tfrac{\pi}{2},\tfrac{\pi}{2}),\forall\omega\in[0,\infty]$ (see [14] for a detailed derivation). Likewise, since $H_{2}(s)$ is passive, it follows that $H_{2}(s)$ is semi-stable with poles at $\mathrm{j}\Omega$ on the imaginary axis and frequency-wise semi-sectorial with phases $\left[\underline{\phi}(H_{2}(\mathrm{j}\omega)),\overline{\phi}(H_{2}(\mathrm{j}\omega))\right]\subset[-\tfrac{\pi}{2},\tfrac{\pi}{2}],\forall\omega\in[0,\infty]\backslash\Omega$ (see [14] for a detailed derivation). Therefore, the conditions in 3 are satisfied which implies $H_{1}\#H_{2}\in\mathcal{RH}_{\infty}^{2n\times 2n}$. ∎
To extend the passivity-based stability conditions to more general feedback interconnections with both passive and nonpassive subsystems, we employ the concepts of excess and shortage of passivity [17, 18]. The basic idea is that the excess of passivity of one subsystem can compensate for the passivity deficit in the other subsystem, such that their feedback interconnection remains stable. This can be achieved by performing loop-shifting techniques as shown in Fig. 1, where a transfer matrix $\Gamma(s)\in\mathbb{C}^{n\times n}$ is added as a positive feedforward to $H_{2}(s)$, and as a positive feedback to $H_{1}(s)$. Now the idea is that if the original subsystems $H_{1}(s)$ and $H_{2}(s)$ in Fig. 1(a) are not satisfying the passivity conditions in Theorem 1, the two subsystems ${H}^{\prime}_{2}(s)$ and ${H}^{\prime}_{1}(s)$ in Fig. 1(b) might do so for a suitable $\Gamma(s)$, thus resulting in a stable feedback interconnection ${H}^{\prime}_{1}\#{H}^{\prime}_{2}$ (for details see literature on frequency-dependent passivity indices, e.g., [17, 18]).
@@ 3.1 Small-Signal Network Dynamics
$\begin{bmatrix}\Delta i_{\mathrm{d},ij}(s)\\ \Delta i_{\mathrm{q},ij}(s)\end{bmatrix}=y_{ij}(s)\left(\begin{bmatrix}\Delta v_{\mathrm{d},i}(s)\\ \Delta v_{\mathrm{q},i}(s)\end{bmatrix}-\begin{bmatrix}\Delta v_{\mathrm{d},j}(s)\\ \Delta v_{\mathrm{q},j}(s)\end{bmatrix}\right),$ (4)
@@
where $b_{ij}=\tfrac{1}{l_{ij}}$ is the line susceptance, and $\rho_{ij}=\tfrac{r_{ij}}{l_{ij}}$ is the resistance-inductance ratio of the line $ij$ which is assumed to be small and uniform (i.e., $\rho_{ij}=\rho\ll 1,\,\forall\,i,j\in\{1,...,n\}$) throughout the dominantly inductive transmission network. Notice that $b_{ij}=0$ and $\rho_{ij}=0$ if there is no line.
@@
$\begin{split}|v|_{i}(t)=\sqrt{v_{\mathrm{d},i}(t)^{2}+v_{\mathrm{q},i}(t)^{2}},\quad\delta_{i}(t)=\arctan\left(\tfrac{v_{\mathrm{q},i}(t)}{v_{\mathrm{d},i}(t)}\right).\end{split}$ (8)
@@
By linearizing 8 and 10, transforming them into the frequency domain, and performing some analytical computations (see Appendix 7 for details), we obtain
$\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta p(s)\\ \Delta q(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta p_{1}(s)\\ \Delta q_{1}(s)\\ \vdots\\ \Delta p_{n}(s)\\ \Delta q_{n}(s)\end{bmatrix}}}=\underset{\eqqcolon\,{N}(s)}{\underbrace{\begin{bmatrix}N_{11}(s)&\cdots&N_{1n}(s)\\ \vdots&\ddots&\vdots\\ N_{n1}(s)&\cdots&N_{nn}(s)\end{bmatrix}}}\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta\delta(s)\\ \Delta|v|_{\mathrm{n}}(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta\delta_{1}(s)\\ \Delta{|v|}_{\mathrm{n},1}(s)\\ \vdots\\ \Delta\delta_{n}(s)\\ \Delta{|v|}_{\mathrm{n},n}(s)\end{bmatrix}}}$ (11)
@@
$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\rho+\tfrac{s}{\omega_{0}}\\ -(\rho+\tfrac{s}{\omega_{0}})&1\end{bmatrix}\\ &\quad\quad\quad+\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&\rho\\ \rho&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}-1&-(\rho+\tfrac{s}{\omega_{0}})\\ (\rho+\tfrac{s}{\omega_{0}})&-1\end{bmatrix},\end{split}$ (12)
where $\Delta{|v|}_{\mathrm{n},i}(s)\coloneqq\tfrac{\Delta|v|_{i}(s)}{|v|_{0,i}}$ is the voltage magnitude at bus $i$ normalized at the steady-state $|v|_{0,i}$. Moreover, to derive 12, as in any power system small-signal model, we have assumed a small steady-state angle difference $\delta_{0,i}\approx\delta_{0,j}$, which is standard and justified since thermal limitations for transmission lines preclude large angle differences [20].
@@
appearing in $N_{ii}(s)$ and $N_{ij}(s)$ in 12 is
$\begin{split}M_{1}(s)&=\tfrac{1}{1+\tfrac{s^{2}}{\omega_{0}^{2}}}\begin{bmatrix}1&\tfrac{s}{\omega_{0}}\\ -\tfrac{s}{\omega_{0}}&1\end{bmatrix},\end{split}$ (14)
@@
$\tfrac{||M(\mathrm{j}\omega)-M_{2}(\mathrm{j}\omega)||_{F}}{||M(\mathrm{j}\omega)-M_{1}(\mathrm{j}\omega)||_{F}}=\sqrt{\tfrac{1+\tfrac{\omega^{4}}{\omega_{0}^{4}}-2\tfrac{\omega^{2}}{\omega_{0}^{2}}}{1+\tfrac{\omega^{4}}{\omega_{0}^{4}}+6\tfrac{\omega^{2}}{\omega_{0}^{2}}+\rho^{2}\left(1+\tfrac{\omega^{2}}{\omega_{0}^{2}}\right)}}<1,$ (16)
that is, $M_{2}$ is a strictly better approximation for $\rho>0$. For $\rho\rightarrow 0$, it can be shown that both approximations are consistent in terms of recovering $M$ asymptotically. Using the approximation in 15, we can therefore eliminate the less dominant terms of the off-diagonal matrix elements in 12 as stated below.
@@
The transfer matrix blocks for the dynamic (i.e., $s\neq 0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 11 used to derive the stability certificates in Section 4 are given by:
$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\tfrac{s}{\omega_{0}}\\ -\tfrac{s}{\omega_{0}}&1\end{bmatrix}\\ &\quad\quad\quad+\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}-1&-\tfrac{s}{\omega_{0}}\\ \tfrac{s}{\omega_{0}}&-1\end{bmatrix}.\end{split}$ (17)
By setting $s=0$ in 17, we derive the quasi-stationary network model, which is employed in a similar form in [9]:
@@
The transfer matrix blocks for the quasi-stationary (i.e., $s=0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 11 are given by:
$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{1}{1+\rho^{2}}\begin{bmatrix}|v|_{0,i}|v|_{0,j}&0\\ 0&2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&-1\end{bmatrix}.\end{split}$ (18)
To further simplify 18, we set $|v|_{0,i}=|v|_{0,j}=|v|_{0}$, yielding the zero-power flow network model similar to [10]:
@@
The transfer matrix blocks for the zero-power flow (i.e., $s=0$, $|v|_{0,i}=|v|_{0,j}=|v|_{0}$) small-signal network model in 11 are given by:
$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0}^{2}}{1+\rho^{2}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0}^{2}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&-1\end{bmatrix}.\end{split}$ (19)
@@
Remark 2.
For $\rho=0$, the network model in [9] corresponds to 18 in Definition 5, while the model in [10] aligns with 19 in Definition 6. Moreover, other works [8, 7] focus solely on SISO frequency dynamics. Hence, our dynamic network model in 17 offers the most detailed representation in literature which is compatible with theoretical stability certificates.

### v6

@@ 1 Introduction
The transition to future power systems is characterized by a substantial increase in the share of power electronics (PE)-based generation devices [1]. This shift introduces significant changes in system dynamics, where the interactions between PE-based generation and the remainder of the power network are not fully understood yet [2, 3, 4]. In particular, the faster dynamics of the control loops and filters in PE-based devices, compared to conventional generators, may induce unforeseen phenomena (e.g., overvoltages[3], subsynchronous oscillations[4]), thereby posing significant challenges to system stability. Consequently, there is a pressing need for a deeper understanding of how PE-based devices interact with one another or with the grid and for the development of stability frameworks to ensure stable operation.
@@
Our contribution improves significantly over prior works on stability certification of GFM converter systems in the small-signal regime. Unlike [8, 9], which are limited to single-input single-output (SISO) dynamics, our approach accommodates multiple-input multiple-output (MIMO) dynamics coupling frequency and voltage, thus providing a more comprehensive analysis of the system behavior. Additionally, we incorporate the network dynamics, extending beyond the quasi-stationary or zero-power flow approximations in [8, 10, 11]. Our decentralized stability conditions enable the use of heterogeneous GFM models, improving over the homogeneity assumptions in [9, 12]. Moreover, unlike [13], our approach does not require detailed knowledge of the network configuration. Collectively, these attributes allow our framework to provide, to the best of our knowledge, the most general, explicitly parametric, and practically applicable small-signal stability certificates for interconnected GFM devices in the literature. Since our stability conditions are readily actionable for control tuning and grid-code design, they can ensure a stable and scalable integration of PE-based generation in future power systems.
@@ 2.1 Preliminaries
Let $\mathbb{R}$ denote the set of real numbers, $\mathbb{N}$ the set of positive integers, and $\mathbb{C}$ the set of complex numbers with imaginary unit $\mathrm{j}$. We use $I_{n}$ to denote the $n$-by-$n$ identity matrix (abbreviated as $I$ when the dimensions can be inferred from the context). We use $\otimes$ to denote the Kronecker product. We use $A=\mathrm{diag}(A_{1},A_{2},...,A_{k}),\,k\in\mathbb{Z}$ to denote the block-diagonal matrix with blocks $A_{1},A_{2},...,A_{k}$. The Euclidean norm of a matrix $A$ is defined as $A||_{F}=\sqrt{\textstyle\sum_{i,j}|a_{ij}|^{2}}$, where $a_{ij}$ is the entry in the $i$th row and $j$th column of $A$.
@@
To establish our main stability results in Section 4, we require the following lemma which follows from the Gershgorin’s Circle Theorem [14, Thm. 6.1.10]:
@@
[…]
@@ 2.2 Transfer Functions & Stability
Let $u(t)\in\mathbb{R}^{n}$ be the input and $y(t)\in\mathbb{R}^{n}$ the output signal of a MIMO LTI system. The $n\times n$ transfer matrix $H(s)$ describes the input-output system response in the frequency domain as $y(s)=H(s)u(s)$, where $y(s)$ and $u(s)$ are the Laplace transformations of the output and input, respectively.
@@
$\hskip-7.11317pt\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}\hskip-3.41432pt=\hskip-3.41432pt\underset{\eqqcolon\,H_{1}\#H_{2}(s)}{\underbrace{\begin{bmatrix}(I\hskip-1.70717pt+\hskip-1.70717ptH_{1}H_{2})^{-1}H_{1}&-(I\hskip-1.70717pt+\hskip-1.70717ptH_{1}H_{2})^{-1}H_{1}H_{2}\\ H_{2}(I\hskip-1.70717pt+\hskip-1.70717ptH_{1}H_{2})^{-1}H_{1}&H_{2}(I\hskip-1.70717pt+\hskip-1.70717ptH_{1}H_{2})^{-1}\end{bmatrix}}}\hskip-4.2679pt\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}\hskip-4.55244pt$ (2)
@@
If there are no RHP pole-zero cancellations between $H_{1}(s)$ and $H_{2}(s)$, then stability of one closed-loop transfer matrix implies stability of the others [16, Thm 4.5]:
@@ 2.3 Small-Phase & Passivity Theory
For LTI systems, the property of passivity is equivalent to the property of positive realness [17, 18, 19].
Definition 2 (Passivity).
A $n\times n$ real rational proper transfer matrix $H(s)$ is said to be passive if
@@
    any purely imaginary pole $\mathrm{j}\omega$ of any element of $H(s)$ is a simple pole and $\text{lim}_{s\rightarrow\mathrm{j}\omega}(s\hskip-1.99168pt-\hskip-1.99168pt\mathrm{j}\omega)H(s)$ is positive semi-definite Hermitian.
Definition 3 (Strict Passivity).
A $n\times n$ real rational proper transfer matrix $H(s)$ is said to be strictly passive if
@@
    $H(\mathrm{j}\omega)+H^{\star}(\mathrm{j}\omega)\succ 0$ for any $\omega\in(-\infty,\infty)$.
We now present a variant of the passivity theorem [19] for LTI systems as follows:
@@
Consider two LTI systems $H_{1}(s)$ and $H_{2}(s)$ in negative feedback configuration, as in Fig. 1(a). The feedback system is internally feedback stable if
  • $H_{1}(s)$ is strictly passive,
  • $H_{2}(s)$ is passive, and
  • $\bar{\sigma}(H_{1}(\mathrm{j}\infty))\bar{\sigma}(H_{2}(\mathrm{j}\infty))<1$.
Several variants of the above theorem exist, often for the SISO case, with subtle distinctions arising particularly in the special case $\omega=\infty$. To avoid ambiguity, we provide a self-contained proof below, relying on the Mixed Gain–Phase Theorem with Cutoff Frequency [20].
Lemma 3 (Mixed Gain-Phase Theorem with Cutoff Frequency).
Let $\omega_{\mathrm{c}}\in(0,\infty)$, $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ be frequency-wise sectorial, and $H_{2}(s)$ be semi-stable frequency-wise semi-sectorial over $(-\omega_{\mathrm{c}},\omega_{\mathrm{c}})$ with $\mathrm{j}\Omega$ being the set of poles on the imaginary axis satisfying $\text{max}_{\omega\in\Omega}|\omega|<\omega_{\mathrm{c}}$. Then, $H_{1}\#H_{2}$ in Fig. 1(a) is internally feedback stable if
  1. (i)
    for each $\omega\in[0,\omega_{\mathrm{c}})\backslash\Omega$, it holds
    $\begin{split}&\overline{\phi}(H_{1}(\mathrm{j}\omega))+\overline{\phi}(H_{2}(\mathrm{j}\omega))<\pi\quad\text{and}\\ &\underline{\phi}(H_{1}(\mathrm{j}\omega))+\underline{\phi}(H_{2}(\mathrm{j}\omega))>-\pi,\end{split}$ (3)
  2. (ii)
    and for each $\omega\in[\omega_{\mathrm{c}},\infty]$, it holds
    $\bar{\sigma}(H_{1}(\mathrm{j\omega}))\bar{\sigma}(H_{2}(\mathrm{j\omega}))<1.$ (4)
@@
We establish Theorem 1 by invoking Lemma 3 with $\omega_{\mathrm{c}}=\infty$. Since Lemma 3 holds for arbitrarily large $\omega_{\mathrm{c}}$, continuity ensures its validity also at $\omega_{\mathrm{c}}=\infty$. Accordingly, we proceed as follows. Because $H_{1}(s)$ is strictly passive, it follows that $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ is frequency-wise sectorial with phases $\left[\underline{\phi}(H_{1}(\mathrm{j}\omega)),\overline{\phi}(H_{1}(\mathrm{j}\omega))\right]\subset(-\tfrac{\pi}{2},\tfrac{\pi}{2}),\forall\omega\in[0,\infty)$. Likewise, since $H_{2}(s)$ is passive, it follows that $H_{2}(s)$ is semi-stable with poles at $\mathrm{j}\Omega$ on the imaginary axis and frequency-wise semi-sectorial with phases $\left[\underline{\phi}(H_{2}(\mathrm{j}\omega)),\overline{\phi}(H_{2}(\mathrm{j}\omega))\right]\subset[-\tfrac{\pi}{2},\tfrac{\pi}{2}],\forall\omega\in[0,\infty]\backslash\Omega$. Hence, the conditions in 3 in Lemma 3 are satisfied. For $\omega=\infty$, it becomes immediate how 4 in Lemma 3 is satisfied. This implies $H_{1}\#H_{2}\in\mathcal{RH}_{\infty}^{2n\times 2n}$. ∎
To extend the passivity-based stability conditions to more general feedback interconnections with both passive and nonpassive subsystems, we employ the concepts of excess and shortage of passivity [18, 19]. The basic idea is that the excess of passivity of one subsystem can compensate for the passivity deficit in the other subsystem, such that their feedback interconnection remains stable. This can be achieved by performing loop-shifting techniques as shown in Fig. 1, where a transfer matrix $\Gamma(s)\in\mathbb{C}^{n\times n}$ is added as a positive feedforward to $H_{2}(s)$, and as a positive feedback to $H_{1}(s)$. Now the idea is that if the original subsystems $H_{1}(s)$ and $H_{2}(s)$ in Fig. 1(a) are not satisfying the passivity conditions in Theorem 1, the two subsystems ${H}^{\prime}_{2}(s)$ and ${H}^{\prime}_{1}(s)$ in Fig. 1(b) might do so for a suitable $\Gamma(s)$, thus resulting in a stable feedback interconnection ${H}^{\prime}_{1}\#{H}^{\prime}_{2}$. Notice that this dynamic loop-shifting approach based on the dynamic transfer function $\Gamma(s)$ effectively corresponds to the concept of frequency-dependent (dynamic) passivity indices[18, 19], which, are generally less conservative than static passivity indices.
@@
Remark 2.
Previous dynamic loop-shifting techniques based on the transfer matrix $\Gamma(s)$ are purely mathematical tools for establishing passivity properties and analytically proving closed-loop stability. They do not involve any hardware changes or modifications to the physical control system.
@@ 3.1 Small-Signal Network Dynamics
[…]
@@
where $b_{ij}=\tfrac{1}{l_{ij}}$ is the line susceptance, and $\rho_{ij}=\tfrac{r_{ij}}{l_{ij}}$ is the resistance-inductance ratio of the line $ij$ which is assumed to be small and uniform (i.e., $\rho_{ij}=\rho\ll 1,\,\forall\,i,j\in\{1,...,n\}$) throughout the dominantly inductive transmission network. If there is no line, we consider $b_{ij}=0$ and $\rho_{ij}=0$. As is standard in small-signal analysis, the network model assumes fixed RL-line parameters and a constant $\omega_{0}$ in the applied $\mathrm{dq}$ transformation, and is therefore valid only in the small-signal regime around the nominal operating frequency.
@@
$\begin{split}\hskip-5.69054pt|v|_{i}(t)\hskip-1.99168pt=\hskip-1.99168pt\sqrt{\hskip-1.99168ptv_{\mathrm{d},i}(t)^{2}\hskip-1.99168pt+\hskip-1.99168ptv_{\mathrm{q},i}(t)^{2}},\quad\delta_{i}(t)\hskip-1.99168pt=\arctan\hskip-1.99168pt\left(\tfrac{v_{\mathrm{q},i}(t)}{v_{\mathrm{d},i}(t)}\right)\hskip-1.42262pt.\hskip-5.69054pt\end{split}$ (9)
@@
By linearizing 9 and 11, transforming them into the frequency domain, and performing some analytical computations (see Appendix 7 for details), we obtain
$\hskip-2.84526pt\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta p(s)\\ \Delta q(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta p_{1}(s)\\ \Delta q_{1}(s)\\ \vdots\\ \Delta p_{n}(s)\\ \Delta q_{n}(s)\end{bmatrix}}}\hskip-1.42262pt=\hskip-1.42262pt\underset{\eqqcolon\,{N}(s)}{\underbrace{\begin{bmatrix}N_{11}(s)&\cdots&N_{1n}(s)\\ \vdots&\ddots&\vdots\\ N_{n1}(s)&\cdots&N_{nn}(s)\end{bmatrix}}}\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta\delta(s)\\ \Delta|v|_{\mathrm{n}}(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta\delta_{1}(s)\\ \Delta{|v|}_{\mathrm{n},1}(s)\\ \vdots\\ \Delta\delta_{n}(s)\\ \Delta{|v|}_{\mathrm{n},n}(s)\end{bmatrix}}}\hskip-2.84526pt$ (12)
@@
[…]
where $\Delta{|v|}_{\mathrm{n},i}(s)\coloneqq\tfrac{\Delta|v|_{i}(s)}{|v|_{0,i}}$ is the voltage magnitude at bus $i$ normalized at the steady-state $|v|_{0,i}$. Moreover, to derive 13, as in any power system small-signal model, we have assumed a small steady-state angle difference $\delta_{0,i}\approx\delta_{0,j}$, which is standard and justified since thermal limitations for transmission lines preclude large angle differences [22].
@@
appearing in $N_{ii}(s)$ and $N_{ij}(s)$ in 13 is
[…]
@@
[…]
that is, $M_{2}$ is a strictly better approximation for $\rho>0$. For $\rho\rightarrow 0$, it can be shown that both approximations are consistent in terms of recovering $M$ asymptotically. Using the approximation 16, we can eliminate the less dominant terms of the off-diagonal matrix elements in 13 as stated below, which allows to exploit symmetry properties of the network dynamics in the stability proof in Section 4.3. However, since 16 is valid only for $\rho\ll 1$, extending our results to more general R/X ratios remains a topic for future work.
@@
The transfer matrix blocks for the dynamic (i.e., $s\neq 0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 12 used to derive the stability certificates in Section 4 are given by:
[…]
By setting $s=0$ in 18, we derive the quasi-stationary network model, which is employed in a similar form in [10]:
@@
The transfer matrix blocks for the quasi-stationary (i.e., $s=0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 12 are given by:
[…]
To further simplify 19, we set $|v|_{0,i}=|v|_{0,j}=|v|_{0}$, yielding the zero-power flow network model similar to [11]:
@@
The transfer matrix blocks for the zero-power flow (i.e., $s=0$, $|v|_{0,i}\hskip-1.42262pt=\hskip-1.42262pt|v|_{0,j}\hskip-1.42262pt=\hskip-1.42262pt|v|_{0}$) small-signal network model in 12 are given by:
[…]
Remark 3.
For $\rho=0$, the network model in [10] corresponds to 19 in Definition 5, while the model in [11] aligns with 20 in Definition 6. Moreover, other works [9, 8] focus solely on SISO frequency dynamics. Hence, our dynamic network model in 18 offers the most detailed representation in literature which is compatible with theoretical stability certificates.